# Combining Evidence Across Filtrations

### v4

@@ Appendix I Cooperative skepticism: A game-theoretic protocol for combining evidence
While we state and prove our main results in the language of measure-theoretic probability, we may also interpret them in the language of game-theoretic probability and statistics (Shafer and Vovk,, 2005, 2019; Ramdas et al.,, 2023). In this view, a test martingale represents the wealth of a gambler who tries to discredit a forecaster by betting against their proposed forecasts (the null hypothesis) over a sequence of outcomes. Specifically, in the testing-by-betting framework (Shafer,, 2021), the wealth of a gambler (“skeptic”) who plays against a point null $P$ is a test (super)martingale for $P$, such that large wealth for the skeptic indicates evidence against $P$. For a composite null $\mathcal{P}$, the gambler is evaluated by the minimum wealth of all skeptics, each of whom plays against each member of the null $P\in\mathcal{P}$. An e-process for $\mathcal{P}$, a composite generalization of a test martingale, corresponds to this minimum wealth against each $P\in\mathcal{P}$ (Ramdas et al.,, 2022). Definition (1) then translates to saying that an e-process can only grow large if the data discredits every possible null scenario.
Following Shafer and Vovk, (2019), we can formalize the game each skeptic plays against a null forecaster $P=(P_{t})_{t\geq 0}$ (the distribution is over the entire data sequence). At each round $t=1,2,\dotsc$, the forecaster proposes a probability distribution $P_{t}$ on the eventual outcome $Y_{t}$, and then a skeptic chooses a bet $S_{t}:\mathcal{Y}\to\mathbb{R}_{\geq 0}$ that determines the multiplicative increment on their wealth depending on the outcome. The game requires that this bet is expected to be small under the forecaster’s proposal:
@@
Based on this framework, we can introduce a game-theoretic protocol that mirrors the setup for Corollary 2. The setup corresponds to having two skeptics that are playing against a common forecaster (the shared null), and their wealth processes correspond to their respective e-processes. In the general setup for this work, each skeptic operates with their own information set, which we will make explicit. For the sake of exposition, we assume that the two e-processes are simply test supermartingales for a point null $P=(P_{t})_{t\geq 0}$ over a data sequence $(Y_{t})_{t\geq 0}$, but the framework can be generalized to e-processes by playing many parallel games for each $P\in\mathcal{P}$ and taking the minimum wealth for each skeptic (Ramdas et al.,, 2022).
Inspired by Dawid et al., 2011b ’s competitive skepticism protocol, we refer to the following protocol as cooperative skepticism to highlight the shared effort among the two skeptics.
@@
We close with the remark that, while Protocol 1 describes a game-theoretic interpretation of Theorem 2 and Corollary 2, it is not itself a result in game-theoretic probability like Dawid et al., 2011b ’s. Dawid et al., 2011b defines a perfect-information protocol, with which they formally prove the equivalence between (11) and (26) to characterize adjusters. On the other hand, Protocol 1 is not a perfect-information game, and it remains unclear whether Theorem 2 can be proven solely in the game-theoretic framework (without measure-theoretic language).
@@
Statistical methods often leverage external randomization to improve their power, and generally, these methods can be readily incorporated into the $\mathfrak{e}$-lifting procedure. For instance, the conformal test martingale in Example 1 uses randomization to break ties (see Section B for details); Ramdas and Manole, (2024, section 10.2) discuss many other known examples. In this section, we explore an external randomization method that may improve the $\mathfrak{e}$-lifting procedure itself.
The main motivation stems from two recent works that demonstrate how utilizing an independent random variable $U\sim\mathsf{Unif}[0,1]$ (“$U$-randomization”) can improve the statistical power. First, in a non-sequential setup, Ignatiadis et al., (2024) remark that if $e$ is an e-value and $U$ is independent of $e$, then $\tilde{p}=(U/e)\wedge 1$ is a valid, $U$-randomized p-value. In fact, the $U$-randomized e-to-p calibrator $e\mapsto(U/e)\wedge 1$ dominates the e-to-p calibrator $e\mapsto(1/e)\wedge 1$ of (8), which is the only admissible deterministic e-to-p calibrator (Vovk and Wang,, 2021, ppn. 2.2). Second, in a sequential setup, Ramdas and Manole, (2024) show that we can analogously incorporate $U$-randomization at a stopping time, increasing the statistical power of the stopped process. (More generally, both of these results, as well as all main results in this section, still hold if $U$ is “stochastically larger” than $\mathsf{Unif}[0,1]$, or $P(U\leq u)\leq u$ for all $u\in[0,1]$. For the sake of exposition, we will let $U$ denote an independent $\mathsf{Unif}[0,1]$ random variable in the below.)
In the following, we first show that a “lift-then-randomize” strategy, inspired by Ramdas and Manole, (2024)’s $U$-randomization method, is naturally applicable to our setup: first apply the $\mathfrak{e}$-lifting procedure (Theorem 2), and then apply $U$-randomization at an arbitrary data-dependent stopping time. Next, we describe an analogous “randomize-then-lift” strategy, corresponding to Ignatiadis et al., (2024)’s $U$-randomized e-to-p calibration method for the non-sequential setup, and show empirically that the strategy does not translate to a valid $\mathfrak{e}$-lifting procedure in general.
@@
Consider an e-process $\mathfrak{e}=(\mathfrak{e}_{t})_{t\geq 0}$ for $\mathcal{P}$ in $\mathbb{G}\subseteq\mathbb{F}$. Given an adjuster $\mathsf{A}$, Theorem 2 says that, for any $\mathbb{F}$-stopping time $\tau$, we have $\mathbb{E}_{P}[\mathsf{A}(\mathfrak{e}_{\tau}^{*})]\leq 1$ for any $P\in\mathcal{P}$. Then, by the uniformly-randomized Markov’s inequality (UMI; Ramdas and Manole,, 2024, theorem 1.2), we have
@@
If we wanted to combine $\tilde{p}_{\tau}^{\mathsf{ltr}}$ with a stopped e-process in $\mathbb{F}$, then we would need to apply an extra p-to-e calibration before being able to average the two. Alternatively, if we wanted to combine the p-value with a stopped p-process in $\mathbb{F}$, then we can use an admissible p-merging function (Vovk et al., 2022b, ) to obtain a combined p-value. Neither of these combination methods, however, will result in a p-process (or an e-process).
@@
Typically, a p-process with stopping time validity can further yield a time-uniform inference procedure that can be monitored continuously due to the equivalence lemma (Section 3.1). However, because the lift-then-randomize procedure of Proposition 2 only yields a stopped p-value and not a p-process, we cannot directly obtain an inference procedure that can be monitored continuously over time. Fortunately, it follows straightforwardly from Ramdas and Manole, (2024)’s randomized Ville’s inequality that, when it comes to deriving a level-$\alpha$ sequential test, we can still monitor the original e-process $\mathfrak{e}$ at the threshold $1/\alpha$ before reaching a stopping time, at which we can use the $U$-randomized p-value.
@@
where $U\sim\mathsf{Unif}[0,1]$ is independent of $\mathbb{G}$ (and thus of $\mathfrak{e}$ and $\tau$). The inequality yields a level-$\alpha$ sequential inference procedure for $\mathcal{P}$ in $\mathbb{G}$: (a) collect data and monitor the e-process in $\mathbb{G}$; (b) if $\mathfrak{e}_{t}$ ever crosses $1/\alpha$, stop and reject the null; (c) otherwise, stop at any $\mathbb{G}$-stopping time $\tau$, draw an independent random variable $U\sim\mathsf{Unif}[0,1]$, and reject the null if $\mathfrak{e}_{\tau}\geq U/\alpha$. Because the threshold $U/\alpha$ is almost surely smaller than the level $1/\alpha$, it gives a better chance of rejecting the null under the alternative at a more lenient threshold than the usual $1/\alpha$. Note that we may only apply randomization at the stopping time $\tau$, as opposed to at all times $t$; see Ramdas and Manole, (2024, section 4) for further discussion.
@@
The second bound is the constant $\tilde{\alpha}=\alpha+\alpha\log(1/\alpha)$. If $\alpha=0.05$, then $\tilde{\alpha}$ is approximately $0.200$, which is substantially larger than $\alpha$. To control the type I error at the level $\tilde{\alpha}=0.05$, we need to use a more stringent level of $\alpha\approx 0.0087$ for $\tilde{\mathfrak{p}}_{\tau^{\mathbb{F}}}^{\mathsf{rtl}}$, but then this may offset gains in the statistical power when compared with the non-randomized $\mathfrak{p}$-lifting procedure (Theorem 1). The proof of this bound (shown below) only uses Markov’s inequality and the equivalence lemma (Ramdas et al.,, 2020; Howard et al.,, 2021). It remains to be seen whether it is possible to tighten this bound for arbitrary $\mathbb{F}$-stopping time $\tau^{\mathbb{F}}$, although the tightness of the first bound (for a specific $\tau^{\mathbb{F}}$) indicates that this bound will not be as small as $\alpha$ in general.
@@
The first bound, $\alpha\mathbb{E}\left[{\mathfrak{e}_{\tau^{\mathbb{F}}}}\right]$, follows directly from the UMI (Ramdas and Manole,, 2024). The second bound can be proved using Markov’s inequality and the equivalence lemma. Compared to the proof of the lifting lemma (Lemma 1), it involves extra steps to handle the external random variable $U$, which makes the event $\{\mathfrak{e}_{\tau}\geq U/\alpha\}$ non-adapted to $\mathbb{G}$ or $\mathbb{F}$. Because $\{\mathfrak{e}_{\tau}\geq U/\alpha\}$ is not adapted to $\mathbb{G}$ or $\mathbb{F}$, the usual equivalence between random time validity and stopping time validity (in $\mathbb{G}$ or $\mathbb{F}$) no longer holds.
@@
By Ramdas et al., (2020, lemma 2), specifically the (iii) $\Rightarrow$ (ii) direction, the statement then holds for all random times $T$ (possibly infinite): for any fixed $\epsilon\in(\alpha,1)$,

### v5

@@ Appendix I Cooperative skepticism: A game-theoretic protocol for combining evidence
While we state and prove our main results in the language of measure-theoretic probability, we may also interpret them in the language of game-theoretic probability and statistics (Shafer and Vovk, 2005, 2019; Ramdas et al., 2023). In this view, a test martingale represents the wealth of a gambler who tries to discredit a forecaster by betting against their proposed forecasts (the null hypothesis) over a sequence of outcomes. Specifically, in the testing-by-betting framework (Shafer, 2021), the wealth of a gambler (“skeptic”) who plays against a point null $P$ is a test (super)martingale for $P$, such that large wealth for the skeptic indicates evidence against $P$. For a composite null $\mathcal{P}$, the gambler is evaluated by the minimum wealth of all skeptics, each of whom plays against each member of the null $P\in\mathcal{P}$. An e-process for $\mathcal{P}$, a composite generalization of a test martingale, corresponds to this minimum wealth against each $P\in\mathcal{P}$ (Ramdas et al., 2022). Definition (1) then translates to saying that an e-process can only grow large if the data discredits every possible null scenario.
Following Shafer and Vovk (2019), we can formalize the game each skeptic plays against a null forecaster $P=(P_{t})_{t\geq 0}$ (the distribution is over the entire data sequence). At each round $t=1,2,\dotsc$, the forecaster proposes a probability distribution $P_{t}$ on the eventual outcome $Y_{t}$, and then a skeptic chooses a bet $S_{t}:\mathcal{Y}\to\mathbb{R}_{\geq 0}$ that determines the multiplicative increment on their wealth depending on the outcome. The game requires that this bet is expected to be small under the forecaster’s proposal:
@@
Based on this framework, we can introduce a game-theoretic protocol that mirrors the setup for Corollary 2. The setup corresponds to having two skeptics that are playing against a common forecaster (the shared null), and their wealth processes correspond to their respective e-processes. In the general setup for this work, each skeptic operates with their own information set, which we will make explicit. For the sake of exposition, we assume that the two e-processes are simply test supermartingales for a point null $P=(P_{t})_{t\geq 0}$ over a data sequence $(Y_{t})_{t\geq 0}$, but the framework can be generalized to e-processes by playing many parallel games for each $P\in\mathcal{P}$ and taking the minimum wealth for each skeptic (Ramdas et al., 2022).
Inspired by Dawid et al. (2011b)’s competitive skepticism protocol, we refer to the following protocol as cooperative skepticism to highlight the shared effort among the two skeptics.
@@
We close with the remark that, while Protocol 1 describes a game-theoretic interpretation of Theorem 2 and Corollary 2, it is not itself a result in game-theoretic probability like Dawid et al. (2011b)’s. Dawid et al. (2011b) defines a perfect-information protocol, with which they formally prove the equivalence between (11) and (26) to characterize adjusters. On the other hand, Protocol 1 is not a perfect-information game, and it remains unclear whether Theorem 2 can be proven solely in the game-theoretic framework (without measure-theoretic language).
@@
Statistical methods often leverage external randomization to improve their power, and generally, these methods can be readily incorporated into the $\mathfrak{e}$-lifting procedure. For instance, the conformal test martingale in Example 1 uses randomization to break ties (see Section B for details); Ramdas and Manole (2026, section 10.2) discuss many other known examples. In this section, we explore an external randomization method that may improve the $\mathfrak{e}$-lifting procedure itself.
The main motivation stems from two recent works that demonstrate how utilizing an independent random variable $U\sim\mathsf{Unif}[0,1]$ (“$U$-randomization”) can improve the statistical power. First, in a non-sequential setup, Ignatiadis et al. (2024) remark that if $e$ is an e-value and $U$ is independent of $e$, then $\tilde{p}=(U/e)\wedge 1$ is a valid, $U$-randomized p-value. In fact, the $U$-randomized e-to-p calibrator $e\mapsto(U/e)\wedge 1$ dominates the e-to-p calibrator $e\mapsto(1/e)\wedge 1$ of (8), which is the only admissible deterministic e-to-p calibrator (Vovk and Wang, 2021, ppn. 2.2). Second, in a sequential setup, Ramdas and Manole (2026) show that we can analogously incorporate $U$-randomization at a stopping time, increasing the statistical power of the stopped process. (More generally, both of these results, as well as all main results in this section, still hold if $U$ is “stochastically larger” than $\mathsf{Unif}[0,1]$, or $P(U\leq u)\leq u$ for all $u\in[0,1]$. For the sake of exposition, we will let $U$ denote an independent $\mathsf{Unif}[0,1]$ random variable in the below.)
In the following, we first show that a “lift-then-randomize” strategy, inspired by Ramdas and Manole (2026)’s $U$-randomization method, is naturally applicable to our setup: first apply the $\mathfrak{e}$-lifting procedure (Theorem 2), and then apply $U$-randomization at an arbitrary data-dependent stopping time. Next, we describe an analogous “randomize-then-lift” strategy, corresponding to Ignatiadis et al. (2024)’s $U$-randomized e-to-p calibration method for the non-sequential setup, and show empirically that the strategy does not translate to a valid $\mathfrak{e}$-lifting procedure in general.
@@
Consider an e-process $\mathfrak{e}=(\mathfrak{e}_{t})_{t\geq 0}$ for $\mathcal{P}$ in $\mathbb{G}\subseteq\mathbb{F}$. Given an adjuster $\mathsf{A}$, Theorem 2 says that, for any $\mathbb{F}$-stopping time $\tau$, we have $\mathbb{E}_{P}[\mathsf{A}(\mathfrak{e}_{\tau}^{*})]\leq 1$ for any $P\in\mathcal{P}$. Then, by the uniformly-randomized Markov’s inequality (UMI; Ramdas and Manole, 2026, theorem 1.2), we have
@@
If we wanted to combine $\tilde{p}_{\tau}^{\mathsf{ltr}}$ with a stopped e-process in $\mathbb{F}$, then we would need to apply an extra p-to-e calibration before being able to average the two. Alternatively, if we wanted to combine the p-value with a stopped p-process in $\mathbb{F}$, then we can use an admissible p-merging function (Vovk et al., 2022) to obtain a combined p-value. Neither of these combination methods, however, will result in a p-process (or an e-process).
@@
Typically, a p-process with stopping time validity can further yield a time-uniform inference procedure that can be monitored continuously due to the equivalence lemma (Section 3.1). However, because the lift-then-randomize procedure of Proposition 2 only yields a stopped p-value and not a p-process, we cannot directly obtain an inference procedure that can be monitored continuously over time. Fortunately, it follows straightforwardly from Ramdas and Manole (2026)’s randomized Ville’s inequality that, when it comes to deriving a level-$\alpha$ sequential test, we can still monitor the original e-process $\mathfrak{e}$ at the threshold $1/\alpha$ before reaching a stopping time, at which we can use the $U$-randomized p-value.
@@
where $U\sim\mathsf{Unif}[0,1]$ is independent of $\mathbb{G}$ (and thus of $\mathfrak{e}$ and $\tau$). The inequality yields a level-$\alpha$ sequential inference procedure for $\mathcal{P}$ in $\mathbb{G}$: (a) collect data and monitor the e-process in $\mathbb{G}$; (b) if $\mathfrak{e}_{t}$ ever crosses $1/\alpha$, stop and reject the null; (c) otherwise, stop at any $\mathbb{G}$-stopping time $\tau$, draw an independent random variable $U\sim\mathsf{Unif}[0,1]$, and reject the null if $\mathfrak{e}_{\tau}\geq U/\alpha$. Because the threshold $U/\alpha$ is almost surely smaller than the level $1/\alpha$, it gives a better chance of rejecting the null under the alternative at a more lenient threshold than the usual $1/\alpha$. Note that we may only apply randomization at the stopping time $\tau$, as opposed to at all times $t$; see Ramdas and Manole (2026, section 4) for further discussion.
@@
The second bound is the constant $\tilde{\alpha}=\alpha+\alpha\log(1/\alpha)$. If $\alpha=0.05$, then $\tilde{\alpha}$ is approximately $0.200$, which is substantially larger than $\alpha$. To control the type I error at the level $\tilde{\alpha}=0.05$, we need to use a more stringent level of $\alpha\approx 0.0087$ for $\tilde{\mathfrak{p}}_{\tau^{\mathbb{F}}}^{\mathsf{rtl}}$, but then this may offset gains in the statistical power when compared with the non-randomized $\mathfrak{p}$-lifting procedure (Theorem 1). The proof of this bound (shown below) only uses Markov’s inequality and the equivalence lemma (Ramdas et al., 2020; Howard et al., 2021). It remains to be seen whether it is possible to tighten this bound for arbitrary $\mathbb{F}$-stopping time $\tau^{\mathbb{F}}$, although the tightness of the first bound (for a specific $\tau^{\mathbb{F}}$) indicates that this bound will not be as small as $\alpha$ in general.
@@
The first bound, $\alpha\mathbb{E}\left[{\mathfrak{e}_{\tau^{\mathbb{F}}}}\right]$, follows directly from the UMI (Ramdas and Manole, 2026). The second bound can be proved using Markov’s inequality and the equivalence lemma. Compared to the proof of the lifting lemma (Lemma 1), it involves extra steps to handle the external random variable $U$, which makes the event $\{\mathfrak{e}_{\tau}\geq U/\alpha\}$ non-adapted to $\mathbb{G}$ or $\mathbb{F}$. Because $\{\mathfrak{e}_{\tau}\geq U/\alpha\}$ is not adapted to $\mathbb{G}$ or $\mathbb{F}$, the usual equivalence between random time validity and stopping time validity (in $\mathbb{G}$ or $\mathbb{F}$) no longer holds.
@@
By Ramdas et al. (2020, lemma 2), specifically the (iii) $\Rightarrow$ (ii) direction, the statement then holds for all random times $T$ (possibly infinite): for any fixed $\epsilon\in(\alpha,1)$,